# Fully Dynamic Euclidean $k$-Means

### v2

@@ 1 Introduction
In Euclidean $k$-means clustering, the input consists of a set of $n$ points $X\subseteq\mathbb{R}^{d}$ in a $d$-dimensional Euclidean space, where each point $x\in X$ has a nonnegative weight $w(x)\geq 0$, and a positive integer $k\in[n]$. We have to compute a set of $k$ centers $S\subseteq\mathbb{R}^{d}$, so as to minimize the objective $\sum_{x\in X}w(x)\cdot\left(\mathrm{dist}(x,S)\right)^{2}$, where $\mathrm{dist}(x,S):=\min_{y\in S}\{\mathrm{dist}(x,y)\}$ and $d(x,y)$ denotes the Euclidean distance between the points $x$ and $y$. Our focus in this paper is on a general setting, where the parameters $k,d$ can not be treated as absolute constants. Also without loss of generality, we assume $d=O(\log n)$ by a standard Johnson-Lindenstrauss transform [JL84].
The study of Euclidean $k$-means dates back to 1950s, starting with the celebrated Lloyd’s algorithm [Llo82]. It is regarded as one of the most fundamental clustering problems, and serves as a basic primitive in unsupervised learning. The problem is APX-hard in Euclidean spaces of dimension $d=\Omega(\log n)$ [ACKS15]. In contrast, for every sufficiently small $\epsilon>0$, it admits a $\text{poly}(1/\epsilon)$-approximation algorithm that runs in $\tilde{O}(n^{1+\epsilon})$ time [lTS24, JJLL25]. [^1] We use $\tilde{O}(\cdot)$ notation to hide polylogarithmic factors in $n$ and the aspect ratio, and polynomial factors in $1/\epsilon$. Using standard coreset based sparsification [DSS24], the runtime can be further reduced to $\tilde{O}(n+k^{1+\epsilon})$. This tradeoff between $\mathrm{poly}(1/\epsilon)$-approximation and $\tilde{O}(n+k^{1+\epsilon})$ runtime is believed to be essentially the best possible.222 When $n=k-1$, $k$-means reduces to closest pair (CP) problem on $n$ points, whose current tradeoff is $\mathrm{poly}(1/\epsilon)$-approximation and $n^{1+\epsilon}$ time via locality sensitive hashing [AI06, AR15]. Whether this tradeoff for CP can be improved, i.e., achieving $\tilde{O}(n)$ time for $O(1)$-approximation, is a long-standing open question and is believed that a lower bound exists [Rub18].
@@
A naive way to solve this problem would be to run the best possible static algorithm from scratch after each update [lTS24, DSS24, JJLL25]; leading to $\text{poly}(1/\epsilon)$-approximation, $\tilde{O}(n+k^{1+\epsilon})$ update time and $O(k)$ recourse. But can we do substantially better? In particular, can we design a dynamic algorithm for this problem with (almost) optimal performance guarantees? We answer this question in the affirmative. Our main result is summarized in the theorem below.
@@
    In recent years, significant attention has been devoted towards designing dynamic algorithms for $k$-means and $k$-median objectives [LV17, CHP+19, HK20, FLNS21, BCLP23, BCG+24, BCF25a]. All these papers, however, focus on solving the problem in general metric spaces, where it is known that any dynamic algorithm has to incur an update time of $\Omega(k)$ [CS04]. In contrast, ours is the first work that designs an algorithm for this problem specifically tailored to the Euclidean setting, and accordingly bypasses the $\Omega(k)$ update time barrier.
@@
    Using our dynamic algorithm as a black-box subroutine, we can obtain a static algorithm for Euclidean $k$-means with $\text{poly}(1/\epsilon)$-approximation ratio and $\tilde{O}(n+k^{1+\epsilon})$ runtime: Indeed, we can first reduce the number of input points to $\tilde{O}(k)$ by computing a coreset [DSS24] in $\tilde{O}(n)$ time, and then insert each of these (weighted) input points into our dynamic algorithm. Our algorithm would handle these $\tilde{O}(k)$ updates in $\tilde{O}(k^{1+\epsilon})$ total time, and in the end it will return a $\text{poly}(1/\epsilon)$-approximate solution to the original set of input points. This matches the current best tradeoff between approximation ratio and runtime in the static setting [lTS24, DSS24, JJLL25], which is also believed to be (almost) optimal (see Footnote 2). Furthermore, it is easy to show that any nontrivial dynamic $k$-means algorithm must necessarily incur $\Omega(1)$ recourse, whereas the recourse guarantee achieved by our algorithm is only $\tilde{O}(1)$. In summary, the approximation ratio, update time and recourse bounds of our algorithm are almost optimal.
Our Techniques. At a high level, our approach towards proving Theorem 1.1 consists of two stages. (i) Take an existing dynamic $k$-means algorithm on general metrics as a starting point. (ii) Speed up the update time of the concerned algorithm in Euclidean spaces, without significantly degrading its approximation ratio and recourse, by designing appropriate geometric data structures. At present, there are only two such dynamic algorithms with close to optimal guarantees (in terms of approximation ratio, update time and recourse) on general metrics [BCG+24, BCF25a].
Among them, the dynamic algorithm developed by [BCG+24] incurs $\Theta(k^{\epsilon})$ recourse. Furthermore, it seems that implementing this framework in Euclidean spaces would inherently incur an exponential in $1/\epsilon$ factor loss in the approximation ratio (see Appendix A for a more detailed discussion). Since we are aiming for $\text{poly}(1/\epsilon)$-approximation ratio and $\tilde{O}(1)$ recourse, we cannot use the algorithm of [BCG+24] as the starting point of our approach.
Instead, we build on top of the dynamic algorithm of [BCF25a], which achieves $O(1)$ approximation ratio, $\tilde{O}(k)$ update time and $\tilde{O}(1)$ recourse on general metrics. To achieve our $k^{\epsilon}$ update time, we must break the $\tilde{O}(k)$ update time barrier of general metrics, and improve the framework of [BCF25a] to utilize the Euclidean structure. In Section 3, we review the basic framework of this algorithm, and outline the major challenges we need to overcome to implement a (modified) version of it in the geometric setting. This particularly requires us to implement several geometric data structures that work against adaptive adversary. Section 4 presents an overview of our geometric data structures. One of our key contributions here is the first consistent hashing scheme [CJK+22] that achieves $\tilde{O}(n^{\epsilon})$ running time per point evaluation with competitive parameters. This new consistent hashing is crucial to ensure that our geometric data structures work against an adaptive adversary. Beyond clustering, our data structures generally address more fundamental tasks including dynamic range queries and maintenance of approximately the nearest assignment, which may be of independent interest. Finally, Section 5 outlines how we modify the framework of [BCF25a] and implement the resulting algorithm using our geometric data structures, leading us to Theorem 1.1. Parts II and III of the paper contains the full version of our algorithm and analysis.
@@
Other Related Work. The Euclidean $k$-means problem has also been extensively studied in the special setting of low-dimensional spaces, where the dimension $d$ is assumed to be a constant. This problem in NP-hard even when $d=2$ [MNV12]. In contrast, it is known to admit a static $(1+\epsilon)$-approximation algorithm that runs in $\tilde{O}\left(2^{(1/\epsilon)^{O(d^{2})}}\cdot n\right)$ time [CFS21].
@@
Finally, we note that an influential line of recent work has focused on two other canonical clustering problems in the dynamic setting: facility location [GHL18, CHP+19, BLP22, BGJ+24] and $k$-center [CGS18, GHL+21, BEF+23, CFG+24, LHG+24, BCF+25b, FS25].
@@ 2 Notations and Preliminaries
We fix the following parameters in the paper: $\epsilon$ is a sufficiently small constant, $n$ is the number of points in the data-set, $d$ is the dimension of the space, and $\Delta$ is the aspect ratio, i.e., the max distance divided by the smallest distance between distinct points. As a typical setup, we assume $\Delta\leq\mathrm{poly}(n)$ is an integer. Without loss of generality, the input points that we consider are from the discrete grid $[\Delta]^{d}:=\{(x_{1},\ldots,x_{d})\mid 1\leq x_{i}\leq\Delta\ \forall 1\leq i\leq d\}$, by a standard rounding and rescaling. Also without loss of generality, we assume $d=O(\log n)$ by a standard Johnson-Lindenstrauss transform [JL84] (see the first paragraph in Section 5.1 for a detailed discussion).
Notations.
We now introduce some key notations that will be used in subsequent sections. For integer $n\geq 1$, denote $[n]:=\{1,\ldots,n\}$. Consider a generic function $f:U\to V$; for $T\subseteq U$ we write $f(T):=\{f(u):u\in T\}$, and for $v\in V$ we write $f^{-1}(v):=\{u\in U\mid f(u)=v\}$. For every two sets $A$ and $B$, we use $A-B$ instead of $A\setminus B$, and whenever $B=\{b\}$ is a singleton, we write $A-b$ instead of $A-\{b\}$ for simplicity. For $x\in[\Delta]^{d},r\geq 0$, let $\textsc{Ball}(x,r):=\{y\in[\Delta]^{d}:\mathrm{dist}(x,y)\leq r\}$ be the metric ball. For $T\subseteq[\Delta]^{d}$, let $\mathrm{diam}(T):=\max_{x,y\in T}\mathrm{dist}(x,y)$ be the diameter of $T$. For every weighted dataset $X\subseteq[\Delta]^{d}$ and center set $S\subseteq[\Delta]^{d}$, the cost of $S$ on $X$ is denoted by $\textsc{cost}(X,S)$ and is defined as $\textsc{cost}(X,S):=\sum_{x\in X}w(x)\cdot\mathrm{dist}(x,S)^{2}$. Accordingly, we define $\textsc{AverageCost}(X,S):=\textsc{cost}(X,S)/w(X)$. When the size of $S$ is one, i.e. $S=\{c\}$, we abbreviate $\textsc{cost}(X,\{c\})$ by $\textsc{cost}(X,c)$ and $\textsc{AverageCost}(X,\{c\})$ by $\textsc{AverageCost}(X,c)$. We denote by $\mathrm{OPT}_{k}(X)$ the cost of the optimum solution of size $k$. For every $W\subseteq[\Delta]^{d}$, we also denote by $\mathrm{OPT}_{k}^{W}(X)$ the cost of the optimum solution of size $k$ which is restricted to open centers only in $W$, i.e. $\mathrm{OPT}_{k}^{W}(X):=\min_{S\subseteq W,|S|=k}\textsc{cost}(X,S)$.
3 Overview of the Dynamic Algorithm of [BCF25a]
In this section, we explain the algorithm of [BCF25a] and its main ingredients that works for general metric spaces. Although their algorithm is primarily designed for $k$-median, it can be seamlessly extended to work for $k$-means as well. See Part III for a thorough analysis.

### v3

@@ 1 Introduction
In Euclidean $k$-means clustering, the input consists of a set of $n$ points $X\subseteq\mathbb{R}^{d}$ in a $d$-dimensional Euclidean space, where each point $x\in X$ has a nonnegative weight $w(x)\geq 0$, and a positive integer $k\in[n]$. We have to compute a set of $k$ centers $S\subseteq\mathbb{R}^{d}$, so as to minimize the objective $\sum_{x\in X}w(x)\cdot\left(\mathrm{dist}(x,S)\right)^{2}$, where $\mathrm{dist}(x,S):=\min_{y\in S}\{\mathrm{dist}(x,y)\}$ and $\mathrm{dist}(x,y)$ denotes the Euclidean distance between the points $x$ and $y$. Our focus in this paper is on a general setting, where the parameters $k,d$ can not be treated as absolute constants. Also without loss of generality, we assume $d=O(\log n)$ by a standard Johnson-Lindenstrauss transform [JL84].
[…]
@@
A naive way to solve this problem would be to run the best possible static algorithm from scratch after each update [laTourS24, DraganovSS24, abs-2504-03513]; leading to $\text{poly}(1/\epsilon)$-approximation, $\tilde{O}(n+k^{1+\epsilon})$ update time and $O(k)$ recourse. But can we do substantially better? In particular, can we design a dynamic algorithm for this problem with (almost) optimal performance guarantees? We answer this question in the affirmative. Our main result is summarized in the theorem below.
@@
    In recent years, significant attention has been devoted towards designing dynamic algorithms for $k$-means and $k$-median objectives [LattanziV17, Cohen-AddadHPSS19, HenzingerK20, FichtenbergerLN21, nips/BhattacharyaCLP23, BhattacharyaCGLP24, BCF24]. All these papers, however, focus on solving the problem in general metric spaces, where it is known that any dynamic algorithm has to incur an update time of $\Omega(k)$ [CzumajS04]. In contrast, ours is the first work that designs an algorithm for this problem specifically tailored to the Euclidean setting, and accordingly bypasses the $\Omega(k)$ update time barrier.
@@
    Using our dynamic algorithm as a black-box subroutine, we can obtain a static algorithm for Euclidean $k$-means with $\text{poly}(1/\epsilon)$-approximation ratio and $\tilde{O}(n+k^{1+\epsilon})$ runtime: Indeed, we can first reduce the number of input points to $\tilde{O}(k)$ by computing a coreset [DraganovSS24] in $\tilde{O}(n)$ time, and then insert each of these (weighted) input points into our dynamic algorithm. Our algorithm would handle these $\tilde{O}(k)$ updates in $\tilde{O}(k^{1+\epsilon})$ total time, and in the end it will return a $\text{poly}(1/\epsilon)$-approximate solution to the original set of input points. This matches the current best tradeoff between approximation ratio and runtime in the static setting [laTourS24, DraganovSS24, abs-2504-03513], which is also believed to be (almost) optimal (see Footnote 2). Furthermore, it is easy to show that any nontrivial dynamic $k$-means algorithm must necessarily incur $\Omega(1)$ recourse, whereas the recourse guarantee achieved by our algorithm is only $\tilde{O}(1)$. In summary, the approximation ratio, update time and recourse bounds of our algorithm are almost optimal.
@@
Other Related Work. The Euclidean $k$-means problem has also been extensively studied in the special setting of low-dimensional spaces, where the dimension $d$ is assumed to be a constant. This problem in NP-hard even when $d=2$ [MahajanNV12]. In contrast, it is known to admit a static $(1+\epsilon)$-approximation algorithm that runs in $\tilde{O}\left(2^{(1/\epsilon)^{O(d^{2})}}\cdot n\right)$ time [Cohen-AddadFS21].
Very recently, for the related Euclidean $k$-median problem, a dynamic $O(\mathrm{poly}(\epsilon^{-1})\log n)$ approximation algorithm with $O(\mathrm{poly}(d)\cdot n^{\epsilon})$ update time has been obtained [GJK+26]. However, because of a fundamental distortion lower bound of the tree embedding technique which this result relies on, it does not generalize to $k$-means and it cannot yield $o(\log n)$ approximation ratio. Hence, this result is largely incomparable to ours.
@@
Finally, we note that an influential line of recent work has focused on two other canonical clustering problems in the dynamic setting: facility location [GoranciHL18, Cohen-AddadHPSS19, BhattacharyaLP22, BhattacharyaGJQ24] and $k$-center [ChanGS18, GoranciHLSS21, BateniEFHJMW23, CrucianiFGNS24, LackiHGJR24, icml25, ForsterS25].
2 Our Techniques
At a high level, our approach towards proving Theorem 1.1 consists of two stages. (i) Take an existing dynamic $k$-means algorithm on general metrics as a starting point. (ii) Speed up the update time of the concerned algorithm in Euclidean spaces, without significantly degrading its approximation ratio and recourse, by designing appropriate geometric data structures. At present, there are only two such dynamic algorithms with close to optimal guarantees (in terms of approximation ratio, update time and recourse) on general metrics [BhattacharyaCGLP24, BCF24].
Among them, the dynamic algorithm developed by [BhattacharyaCGLP24] incurs $\Theta(k^{\epsilon})$ recourse. Furthermore, it seems that implementing this framework in Euclidean spaces would inherently incur an exponential in $1/\epsilon$ factor loss in the approximation ratio (see Appendix A for a more detailed discussion). Since we are aiming for $\text{poly}(1/\epsilon)$-approximation ratio and $\tilde{O}(1)$ recourse, we cannot use the algorithm of [BhattacharyaCGLP24] as the starting point of our approach.
Instead, we build on top of the dynamic algorithm of [BCF24], which achieves $O(1)$ approximation ratio, $\tilde{O}(k)$ update time and $\tilde{O}(1)$ recourse on general metrics. To achieve our $k^{\epsilon}$ update time, we break the $\tilde{O}(k)$ update time barrier of general metrics, and adapt the framework of [BCF24] to utilize the Euclidean structure. Although our algorithm is built on the skeleton of the [BCF24] framework, we face intricate challenges that we highlight below, as well as our key technical contributions that show how to get around these challenges.
2.1 Challenge I: Approximate Neighborhood
All of the main objects used in the algorithm of [BCF24] work with exact balls around points in general metric spaces. On the other hand, in Euclidean spaces, the fast geometric data structures can only guarantee an approximate neighborhood of some radius $r$ around a point $x$ if we use any fast geometric data structure, i.e., there might be some points of distance $(1+\epsilon)r$ to $x$ that are mistakenly included in the neighborhood. As a result, the first challenge here is to extend the framework of [BCF24] and come up with new objects that work even with an approximate neighborhood. Unfortunately, the very first main crucial object in [BCF24] called ‘robust center’ must be reconsidered, and does not work with an approximate neighborhood. This means that after our modification, we need to rebuild the framework of [BCF24] from scratch.
We modify the definition deliberately and provide a more complicated definition of ‘robust center’ such that: 1) it can be handled with the approximate neighborhood oracles, and 2) the rest of the framework of [BCF24] can be extended to work w.r.t. the new definition, without introducing a conceptual barrier inside the analysis. Intuitively, a robust center $x$ must satisfy specific properties w.r.t. the balls around $x$ of geometrically increasing radius $10^{0},10^{1},10^{2},\ldots$ (see Definition 4.1). The new definition of ‘robust center’ that we consider generalizes the previous definition in an abstract way, such that no matter what the approximate neighborhood around a point $x$ is (a subset $B$ of the data-set that contains all points of distance at most $r$ to $x$ and some of the points of distance at most $(1+\epsilon)r$ to $x$), it always works with this neighborhood (as well as any other arbitrary neighborhood). This is a non-trivial modification which is too technical, and we refer the reader to Definition 4.1 and Definition 6.1 for a direct comparison.
2.2 Challenge II: Geometric Data Structures Against Adaptive Adversaries
To break the $O(k)$ time barrier of [BCF24], we need to design efficient data structures that leverage the Euclidean structure and support several fundamental geometric operations required in the algorithm of [BCF24]. In our setting, the data structures must handle both a dynamic dataset $X$, which is updated by an adversary, and a dynamic solution (center set) $S$, which is explicitly maintained and updated by the [BCF24] algorithm. Therefore, another major challenge arises: even if we assume that $X$ is updated by an oblivious adversary (so that the whole dynamic $k$-means algorithm faces an oblivious adversary), the updates to $S$ cannot be considered as oblivious, since the [BCF24] algorithm is randomized and makes subsequent updates to $S$ randomly according to the current outputs/states of the data structures. As a result, it is necessary to provide data structures that work against an adaptive adversary. This challenge arises from adapting any randomized algorithm not specifically [BCF24].
To build such data structures, one of our key contributions is a new construction of a consistent hashing scheme [CJKVY22] that, for the first time, achieves sublinar $\tilde{O}(n^{\epsilon})$ running time per point evaluation with competitive parameters. Based on this new consistent hashing, we design robust data structures that can handle two fundamental tasks, answering range queries and maintaining approximate assignments, with $\tilde{O}(n^{\epsilon})$ update and query time; see Section 5 for a more detailed exposition of our consistent hashing and data structure design.
These data structures play an important role in speeding up the [BCF24] algorithm, especially for the key subroutines discussed in the below Section 2.3, and we believe they may be of independent interest, as they address fundamental tasks beyond clustering.
2.3 Challenge III: New Subroutines
All of the critical subroutines in the [BCF24] algorithm can not be directly implemented with fast geometric data structures in the Euclidean spaces, in particular, the three subroutines: ‘Restricted $k$-Means’, ‘Augmented $k$-Means’, and ‘Robustify’. In the following, we explain what the conceptual barriers are towards directly implementing these subroutines as stated in [BCF24], and provide our brand new algorithms for these subroutines that are implementable in Euclidean spaces. Except for ‘Robustify’, we believe that the other two subroutines are handy, and our new algorithms for these subroutines might be of independent interest to the clustering community.
Restricted $k$-Means. In the restricted $k$-means subroutine, the goal is to remove a given number of centers $r\leq k$ from the solution set $S$ so as to minimize the cost among all possible deletions of $r$ centers out of $S$. The [BCF24] algorithm uses randomized local search for this subroutine, which is designed for general metric spaces and thus the inherent $\Omega(kr)$ time lower bound prevents it from achieving sublinear $o(k)$ update time. Here we take a completely different coreset-based approach: we show that there exists a small $O(r)$-sized center subset of $S$ such that solving (approximately) the restricted $k$-means on this subset already suffices. This coreset-based algorithm is more promising in leveraging the Euclidean structure. Indeed, we reduce the construction of such a coreset to approximate nearest neighbor search (which is supported by our data structures), and by applying an efficient algorithm for Euclidean $k$-means [laTourS24, abs-2504-03513] to this coreset, we achieve a nearly tight running time of $\tilde{O}(r^{1+\epsilon})$ for the restricted $k$-means subroutine. Please see Section 5.4.1 for a more complete explanation of this part.
Augmented $k$-Means. In the augmented $k$-means subroutine, the goal is to add $r$ more centers to the solution set $S$ so as to minimize the cost among all possible additions of $r$ centers to $S$, for a given $r>0$. The algorithm proposed in [BCF24] contracts the centers in $S$ and introduces a new metric space such that any $(r+1)$-means solution in the new space corresponds to a solution for the augmented $k$-means problem. It is obvious that the new metric space is far from being Euclidean, and we can not use our Euclidean data structures in order to solve the $r$-means problem in the new metric space. Hence, we need a new algorithm to solve augmented $k$-means. Our key observation here is that the $k$-means++ algorithm [ArthurV07], which iteratively performs $D^{2}$-sampling to add centers, can in fact be used to yields a good approximation for augmented $k$-means. Notice that our data structures can maintain (implicitly and approximately) the nearest-center assignments efficiently. With this support, this new $D^{2}$-sampling-based algorithm can be readily implemented in sublinear time $\tilde{O}(r\cdot n^{\epsilon})$. Please see Section 5.4.2 for more details.
Robustify. In the robustify subroutine, the goal is to detect all the centers in $S$ that might violate the condition of being a robust center, and then swap these centers with a robust center in their neighborhood. The algorithm that [BCF24] proposes here is simply iterating over all elements of $S$, checking the robustness condition, and swap non-robust centers with robust ones. Obviously, this takes $\Omega(k)$ time, which is not affordable for our purpose. According to our modification of the main objects of [BCF24], in order to have a fast ‘Robustify’ subroutine, the main operation that we have to carry out deep inside the subroutine is to detect the centers $u$ in $S$ whose distances to their closest center in the SuS-u has been changed significantly, i.e., multiplicatively by a constant factor. Due to our modification, for each geometrically increasing radius $10^{i}$, we combine our approximate nearest neighbor oracle with one of the data structures in [BhattacharyaGJQ24] in order to detect such centers. Please see Section 6.4 for a more complete explanation of this part.
@@ 3 Notations and Preliminaries
We fix the following parameters in the paper: $\epsilon$ is a sufficiently small constant, $n$ is the number of points in the data-set, $d$ is the dimension of the space, and $\Delta$ is the aspect ratio, i.e., the maximum distance divided by the smallest distance between distinct points. As a typical setup, we assume $\Delta\leq\mathrm{poly}(n)$ is an integer. Without loss of generality, the input points that we consider are from the discrete grid $\Delta]^{d}:=\{(x_{1},\ldots,x_{d})\mid 1\leq x_{i}\leq\Delta\ \forall 1\leq i\leq d\}$, by a standard rounding and rescaling. Also without loss of generality, we assume $d=O(\log n)$ by a standard Johnson-Lindenstrauss transform [JL84] (see the first paragraph in Section 6.1 for a detailed discussion).
Notations. We now introduce some key notations that will be used in subsequent sections. For integer $n\geq 1$, denote $n]:=\{1,\ldots,n\}$. Consider a generic function $f:U\to V$; for $T\subseteq U$ we write $f(T):=\{f(u):u\in T\}$, and for $v\in V$ we write $f^{-1}(v):=\{u\in U\mid f(u)=v\}$. For every two sets $A$ and $B$, we use ABA-B instead of $A\setminus B$, and whenever $B=\{b\}$ is a singleton, we write AbA-b instead of $A-\{b\}$ for simplicity. For $x\in[\Delta]^{d},r\geq 0$, let $\textsc{Ball}(x,r):=\{y\in[\Delta]^{d}:\mathrm{dist}(x,y)\leq r\}$ be the metric ball. For $T\subseteq[\Delta]^{d}$, let $\mathrm{diam}(T):=\max_{x,y\in T}\mathrm{dist}(x,y)$ be the diameter of $T$. For every weighted dataset $X\subseteq[\Delta]^{d}$ and center set $S\subseteq[\Delta]^{d}$, the cost of $S$ on $X$ is denoted by $\textsc{cost}(X,S)$ and is defined as $\textsc{cost}(X,S):=\sum_{x\in X}w(x)\cdot\mathrm{dist}(x,S)^{2}$. Accordingly, we define $\textsc{AverageCost}(X,S):=\textsc{cost}(X,S)/w(X)$. When the size of $S$ is one, i.e. $S=\{c\}$, we abbreviate $\textsc{cost}(X,\{c\})$ by $\textsc{cost}(X,c)$ and $\textsc{AverageCost}(X,\{c\})$ by $\textsc{AverageCost}(X,c)$. We denote by $\mathrm{OPT}_{k}(X)$ the cost of the optimum solution of size $k$. For every $W\subseteq[\Delta]^{d}$, we also denote by $\mathrm{OPT}_{k}^{W}(X)$ the cost of the optimum solution of size $k$ which is restricted to open centers only in $W$, i.e. $\mathrm{OPT}_{k}^{W}(X):=\min_{S\subseteq W,|S|=k}\textsc{cost}(X,S)$.
Organization. We provide an overview of the algorithm of [BCF24] in Section 4. We provide our geometric data structures in Section 5. In Section 6, we provide our modification of the algorithm of [BCF24]. Finally, in Section 7, we provide the sketch of the analysis of our algorithm. We refer the reader to Part II for a thorough description and analysis of our geometric data structures, and to Part III for a complete description and analysis of all parts of our algorithm.
4 Overview of the Dynamic Algorithm of [BCF24]
In this section, we explain the algorithm of [BCF24] and its main ingredients that works for general metric spaces. Although their algorithm is primarily designed for $k$-median, it can be seamlessly extended to work for $k$-means as well. See Part III for a thorough analysis.